# Relative-error monotonicity testing

### v1

@@ 6.2 Proof of Lemma 23
We finish the proof by proving the two claims:
Proof of 28.
Fix a good $\ell=(P;R;\rho)$ and let $I\mathrel{\mathop{\ordinarycolon}}=\{i\in[L]\mathrel{\mathop{\ordinarycolon}}P_{i}\neq\emptyset\}$. For each $i\in I$, let
@@
Since the outcome $\ell$ is good, we have
@@
    |$\big{|}\{k\in[n]\mathrel{\mathop{\ordinarycolon}}x_{k}=y_{k}=1\}\big{|}>3n/4-100\log n$
@@
    |$\big{|}\{k\in[n]\mathrel{\mathop{\ordinarycolon}}x_{k}=1\cap y_{k}=0\}\big{|}=\|x\|_{1}-\big{|}\{k\in[n]\mathrel{\mathop{\ordinarycolon}}x_{k}=y_{k}=1\}\big{|}\leq 100\log n.$
    For each $i\in[L]$ with any $x\in P_{i}$, we have
@@
  • For each $i\in I$, we use $\rho_{i}$ to denote the value of $\rho_{i}(x)$ shared by all $x\in P_{i}$.
Consider a fixed $T$ in the support of $\mathcal{E}$ such that the probability of $(T,{\bm{H}})$ resulting in outcome $\ell$ is positive when ${\bm{H}}\sim\mathcal{E}_{\text{yes}}$. Then it suffices to show that
$\dfrac{\operatorname{{\bf Pr}}_{{\bm{H}}\sim\mathcal{E}_{\text{no}}}[(T,{\bm{H}})\text{ results in outcome $\ell$}]}{\operatorname{{\bf Pr}}_{{\bm{H}}\sim\mathcal{E}_{\text{yes}}}[(T,{\bm{H}})\text{ results in outcome $\ell$}]}\geq\frac{1}{1+\alpha}.$
By definition of the distributions $\mathcal{E}_{\text{no}}$ and $\mathcal{E}_{\text{yes}}$, this ratio is a product over $i\in I$. For each $i\in I$, if $\rho_{i}=1$, then the factor is
$\frac{\frac{2}{3}\cdot\frac{|A_{i,0}|}{n}+\frac{1}{3}}{\frac{2}{3}\cdot\frac{|A_{i,1}|}{n}}=\frac{(2/3)\cdot|A_{i,0}|+(n/3)}{(2/3)\cdot|A_{i,1}|}.$
Given that $|A_{i,1}|=(3n/4)\pm O(|P_{i}|\log n)$ and $|A_{i,0}|=(n/4)\pm O(|P_{i}|\log n)$ (and $|P_{i}|\log n\leq 2q\log n$ which can be set to be sufficiently small compared to $n$ by setting $C_{0}$ to be sufficiently small),
$\frac{(2/3)\cdot|A_{i,0}|+(n/3)}{(2/3)\cdot|A_{i,1}|}=\frac{(n/2)\pm O(|P_{i}|\log n)}{(n/2)\pm O(|P_{i}|\log n)}=1\pm O\left(\frac{|P_{i}|\log n}{n}\right).$
Similarly when $\rho_{i}=0$, the factor is
$\frac{\frac{2}{3}\cdot\frac{|A_{i,1}|}{n}}{\frac{2}{3}\cdot\frac{|A_{i,0}|}{n}+\frac{1}{3}}=\frac{(2/3)\cdot A_{i,1}}{(2/3)\cdot A_{i,0}+(n/3)}=1\pm O\left(\frac{|P_{i}|\log n}{n}\right).$
As a result, using $\sum_{i\in I}|P_{i}|\leq 2q=2n/(C_{0}\log n)$ and setting $C_{0}$ to be sufficiently small, we have
$\dfrac{\operatorname{{\bf Pr}}_{{\bm{H}}\sim\mathcal{E}_{\text{no}}}[(T,{\bm{H}})\text{ results in outcome $\ell$}]}{\operatorname{{\bf Pr}}_{{\bm{H}}\sim\mathcal{E}_{\text{yes}}}[(T,{\bm{H}})\text{ results in outcome $\ell$}]}$ $\geq\prod_{i\in I}\left(1-O\left(\dfrac{|P_{i}|\log n}{n}\right)\right)\geq\frac{1}{1+\alpha}.$
@@
Proof of 27.
We first handle the first case of bad outcomes: some $i\in[L]$ has $x,y\in P_{i}$ with
|$\big{|}\{k\in[n]\mathrel{\mathop{\ordinarycolon}}x_{k}=y_{k}=1\}\big{|}\leq(3n/4)-100\log n.$
By a union bound over $x,y$ and $i$, this happens with probability at most
$(4/3)^{n}\cdot q^{2}\cdot\left(\dfrac{(3n/4)-100\log n}{n}\right)^{n}=q^{2}\left(1-\frac{(4/3)100\log n}{n}\right)^{n}=o(1).$
Assume the first case does not happen. Define the same $I$ and $A_{i,0},A_{i,1}$ as in the last proof. Using the same inequalities on $|A_{i,0}|$ and $|A_{i,1}|$, we have for every $i\in I$,
$\Big{|}[n]\setminus(A_{i,0}\cup A_{i,1})\Big{|}\leq O(|P_{i}|\log n).$
For the second case to happen, there must be an $i\in I$ such that ${\bm{h}}_{i}(x)=x_{k}$ sampled according to $\mathcal{E}_{\text{yes}}$ has $k\in[n]\setminus(A_{i,0}\cup A_{i,1})$. Thus, taking a union bound over all possible $i\in I$, the probability over ${\bm{H}}\sim\mathcal{E}_{\text{yes}}$ of resulting in such an outcome is at most
$\sum_{i\in I}\left(\frac{2}{3}\cdot\dfrac{\big{|}[n]\setminus(A_{i,0}\cup A_{i,1})\big{|}}{n}\right)\leq\sum_{i\in I}\left(\dfrac{O(|P_{i}|\log n)}{n}\right)\leq\alpha,$
by using $\sum_{i\in I}|P_{i}|\leq 2q=2n/(C_{0}\log n)$ from 25 and setting $C_{0}$ to be sufficiently small. ∎
@@ 7 A two-sided adaptive lower bound
Observations similar to those below the statement of Theorem 19 show that Theorem 29 implies Theorem 3. As it will become clear in Section 7.1, all functions used in our lower bound proof are two-layer functions as well. It follows from 16 that it suffices to show that any randomized, adaptive, $\mathrm{MQ}$-only algorithm for testing two-layer functions with success probability $(99/100)\cdot 2/3$ must make $\tilde{\Omega}((\log n)^{2/3})$ queries. So in the rest of the section, we focus on $\mathrm{MQ}$-only algorithms.
@@ 7.1 Distributions $\mathcal{D}_{\mathrm{yes}}$ and $\mathcal{D}_{\mathrm{no}}$
    Sample ${\bm{H}}=(\bm{h}_{i,j}\mathrel{\mathop{\ordinarycolon}}i\in[L],j\in[M])$ from $\mathcal{E}_{\text{yes}}$, where each ${\bm{h}}_{i,j}\mathrel{\mathop{\ordinarycolon}}\{0,1\}^{n}\rightarrow\{0,1\}$ is (1) with probability $2/3$, a random dictatorship Boolean function, i.e., ${\bm{h}}_{i}(x)=x_{k}$ with $k$ sampled uniformly at random from $[n]$; and (2) with probability $1/3$, the all-00 function.
@@
Note that the number of points on the $43n/$ 4 layer is (${n\choose 3n/4}\geq|f^{-1}(1)|/100$. Thus, the expected size of $\bm{X}$ is at least $|f^{-1}(1)|/10000$, and $|\bm{X}|=\Omega(|f^{-1}(1)|)$ with probability at least 0.0001. ∎

### v2

@@ 5.3 Proof of Lemma 21
5.4 Proofs of Claims
Now we prove 26 and 27. We start with 26.
We divide the proof of 26 into two steps. First we introduce the following distribution $\mathcal{H}_{\text{yes}}^{*}$ for the analysis of the first stage of ALG. To draw a tuple $\bm{U}^{*}=(\bm{S},(\mathbf{P}^{*};\bm{R}^{*};\bm{\rho}^{*}))\sim\mathcal{H}_{\text{yes}}^{*}$, we first draw a function ${\bm{f}}\sim\mathcal{D}_{\mathrm{yes}}$ and a sequence of $q$ points $\bm{S}$ uniformly at random from points that are at or above layer $3n/4$. The triple $(\mathbf{P}^{*};\bm{R}^{*};\bm{\rho}^{*})$ is then set to be $\mathsf{outcome}({\bm{f}},\bm{S})$.
Definition 28.
Let $U^{*}=(S,(P^{*};R^{*};\rho^{*}))$ be a tuple in the support of $\mathcal{H}_{\text{yes}}^{*}$. We say $U^{*}$ is regular if it satisfies the following two conditions: (1) $|P^{*}_{i}|\leq 1$ for all $i\in[L]$, and (2) every two strings in $S$ have at most $5n/8$ many common $1$-entries.
Claim 29.
With probability at least $1-o_{n}(1)$, $\bm{U}^{*}=(\bm{S},(\mathbf{P}^{*};\bm{R}^{*};\bm{\rho}^{*}))\sim\mathcal{H}_{\text{yes}}^{*}$ is regular.
Proof of 29.
Let $\bm{x}^{1},\ldots,\bm{x}^{q}$ be the sequence of samples in $\bm{S}$ (drawn uniformly at random from layer $3n/4$ or above). It follows from Chernoff bound and a union bound that with probability at least $1-o_{n}(1)$, we have for every $1\leq i<j\leq q$, the number of $k\in[n]$ with $x^{i}_{k}=x^{j}_{k}=1$ is at most $5n/8$ (note that the expectation is about $9n/16$). Assuming this is the case for a fixed set of samples $x^{1},\ldots,x^{q}$, it suffices to show that when $\bm{T}_{1},\ldots,\bm{T}_{L}$ are drawn independently and uniformly at random, no $i,j,k$ satisfy $\bm{T}_{k}(x^{i})=\bm{T}_{k}(x^{j})$ with probability at least $1-o_{n}(1)$.
This is because by our assumption on $x^{i}$ and $x^{j}$, the probability of $\bm{T}_{k}(x^{i})=\bm{T}_{k}(x^{j})=1$ for any triple $i,j,k$ is at most $(5/8)^{n}$. Given that $L=(4/3)^{n}$ and $q$ is at most polynomial in $n$, the claim follows from a union bound over all $Lq^{2}$ many triples $i,j,k$. ∎
Now we work on the second stage of ALG. Fix any regular $U^{*}=(S,(P^{*};R^{*};\rho^{*}))$ in the support of $\mathcal{H}_{\text{yes}}^{*}$, and let us write $Q=Q_{U^{*}}$ to denote the set of $q$ queries ALG would query in the second stage when (1) the samples it queries are $S$ in the first stage and (2) the query results of $S$ are determined using $(P^{*};R^{*};\rho^{*})$ as discussed in 23. We write $\mathcal{D}_{\mathrm{yes}}^{\prime}$ to denote $\mathcal{D}_{\mathrm{yes}}$ conditioning on ${\bm{f}}\sim\mathcal{D}_{\mathrm{yes}}$ satisfying $(P^{*};R^{*};\rho^{*})=\mathsf{outcome}({\bm{f}},S)$, and $\mathcal{H}_{\text{yes}}^{\prime}$ to denote the distribution of $\bm{U}=(Q,S,(\mathbf{P};\bm{R};\bm{\rho}))$ by first drawing ${\bm{f}}\sim\mathcal{D}_{\mathrm{yes}}^{\prime}$ and then setting $(\mathbf{P};\bm{R};\bm{\rho})=\mathsf{outcome}({\bm{f}},Q\cup S)$.
We prove the following claim about $\mathcal{H}_{\text{yes}}^{\prime}$ (after fixing any regular $U^{*}$ in the support of $\mathcal{H}_{\text{yes}}^{*}$):
Claim 30.
With probability at least $1-\alpha-o_{n}(1)$, $\bm{U}\sim\mathcal{H}_{\text{yes}}^{\prime}$ is good.
26 follows directly by combining 29 and 30.
Proof of 30.
Fix a regular $U^{*}=(S,(P^{*};R^{*};\rho^{*}))$ in the support of $\mathcal{H}_{\text{yes}}^{*}$ and use it to define $Q,\mathcal{D}_{\mathrm{yes}}^{\prime}$ and $\mathcal{H}_{\text{yes}}^{\prime}$ as above. Let $(Q,S,(\mathbf{P};\bm{R};\bm{\rho}))\sim\mathcal{H}_{\text{yes}}^{\prime}$.
We start by showing that the first bad event, i.e. there exist $x,y\in\mathbf{P}_{i}$ for some $i\in[L]$ satisfying
$\big{|}\{k\in[n]\mathrel{\mathop{\ordinarycolon}}x_{k}=y_{k}=1\}\big{|}\leq(3n/4)-100\log n,$ (6)
happens with probability at most $o_{n}(1)$. After proving this, we condition on it not happening and show that the second bad event, i.e., $\rho_{i}(x)\neq\rho_{i}(y)$ for some $i\in[L]$ and $x,y\in\mathbf{P}_{i}$, happens with probability at most $\alpha$. The claim then follows by a union bound.
For the first bad event, we start with the event that there exists some $i\in[L]$ such that $P_{i}^{*}=\emptyset$ (but $R^{*}_{i}$ needs not to be empty in general) but some $x,y\in\mathbf{P}_{i}$ satisfy Equation 6. We prove this happens with $o_{n}(1)$ probability by a union bound over all $x,y\in Q$ (that satisfy Equation 6) and all $i\in[L]$ with $P_{i}^{*}=\emptyset$. For each such $x,y,i$, we note that $\bm{T}_{i}$ (when we draw ${\bm{f}}\sim\mathcal{D}_{\mathrm{yes}}^{\prime}$) is drawn uniformly at random from all terms $T_{i}$ that satisfy $T_{i}(z)=0$ for all $z\in R_{i}^{*}$. So the probability of this event is
[…]
where $\bm{T}_{i}$ is drawn uniformly at random. It is clear that the denominator is at least $1-o_{n}(1)$ given that all $z\in R_{i}^{*}$ have weight at most $(3n/4)+1$. The numerator, on the other hand, is at most
$\mathop{{\bf Pr}\/}_{\bm{T}_{i}}\big{[}\bm{T}_{i}(x)=\bm{T}_{i}(y)=1\big{]}\leq\left(\dfrac{(3n/4)-100\log n}{n}\right)^{n}=\left(\frac{3}{4}\right)^{n}\left(1-\frac{(4/3)100\log n}{n}\right)^{n}.$
By a union bound over $q^{2}L$ triples of $x,y,i$, we have that this case happens with probability $o_{n}(1)$.
We next handle the event that there exists some $i\in[L]$ with $|P_{i}^{*}|=1$ such that some $x,y\in\mathbf{P}_{i}$ satisfy Equation 6. By triangle inequality, a necessary condition for this event to happen is that there exists some $i\in[L]$ with $|P_{i}^{*}|=1$ (say $P_{i}^{*}=\{x^{*}\}$) such that some $y\in\mathbf{P}_{i}$ satisfies
$\big{|}\{k\in[n]\mathrel{\mathop{\ordinarycolon}}x^{*}_{k}=y_{k}=1\}\big{|}\leq(3n/4)-50\log n.$ (7)
Fix an $i\in[L]$ with $|P_{i}^{*}|=1$ (denoting $P_{i}^{*}=\{x^{*}\}$) and a $y\in Q$ such that Equation 7 holds. We bound the probability of $y\in\mathbf{P}_{i}$ and then apply a union bound over the at most $2q^{2}$ many pairs of $i$ and $y$.
To this end, note that $\bm{T}_{i}$ is drawn uniformly but subject to conditioning on $\bm{T}_{i}(x^{*})=1$ and $\bm{T}_{i}(z)=0$ for all $z\in R^{*}_{i}$. So the probability we are interested in is
$\frac{\operatorname{{\bf Pr}}_{\bm{T}_{i}}[\bm{T}_{i}(x^{*})=\bm{T}_{i}(y)=1\ \text{and}\ \bm{T}_{i}(z)=0\ \text{for all $z\in R_{i}^{*}$}]}{\operatorname{{\bf Pr}}_{\bm{T}_{i}}[\bm{T}_{i}(x^{*})=1\ \text{and}\ \bm{T}_{i}(z)=0\ \text{for all $z\in R_{i}^{*}$}]},$
where $\bm{T}_{i}$ is drawn uniformly at random. Note that the denominator is at least
$(3/4)^{n}\cdot\left(1-|R_{i}^{*}|\cdot\left(\frac{5/8}{3/4}\right)^{n}\right)\geq(1-o_{n}(1))\cdot(3/4)^{n},$
where the $(3/4)^{n}$ is the probability of $\bm{T}_{i}(x^{*})=1$ and the rest is a lower bound for the probability of $\bm{T}_{i}(z)=0$ for all $z\in R^{*}_{i}$ conditioning on $\bm{T}_{i}(x^{*})=1$, using the assumption that $x^{*}$ and $z$ share at most $5n/8$ common $1$-entries. The numerator on the other hand is at most
$\mathop{{\bf Pr}\/}_{\bm{T}_{i}}\big{[}\bm{T}_{i}(x^{*})=\bm{T}_{i}(y)=1\big{]}\leq\left(\dfrac{(3n/4)-50\log n}{n}\right)^{n}=\left(\frac{3}{4}\right)^{n}\left(1-\frac{(4/3)50\log n}{n}\right)^{n}.$
The probability of this case is at most $o_{n}(1)$ by taking the ratio and a union bound over the at most $2q^{2}$ many pairs of $i$ and $y$.
Assume the first bad event does not happen. Fix any such $T$, as well as $P$ and $R$ given $T$. Let $I\mathrel{\mathop{\ordinarycolon}}=\{i\in[L]\mathrel{\mathop{\ordinarycolon}}P_{i}\neq\emptyset\}$ and for each $i\in I$, let
@@ 5.4 Proofs of Claims
Using the assumption that the first bad event does not happen, we have:
@@
    $\big{|}\{k\in[n]\mathrel{\mathop{\ordinarycolon}}x_{k}=y_{k}=1\}\big{|}>3n/4-100\log n$
@@
    $\big{|}\{k\in[n]\mathrel{\mathop{\ordinarycolon}}x_{k}=1\cap y_{k}=0\}\big{|}=\|x\|_{1}-\big{|}\{k\in[n]\mathrel{\mathop{\ordinarycolon}}x_{k}=y_{k}=1\}\big{|}\leq 100\log(n)+1.$
    For any $i\in[L]$ with any $x\in P_{i}$, we have
@@
Now recall that the second bad event is that $\rho_{i}(x)\neq\rho_{i}(y)$ for some $i\in[L]$ and $x,y\in\mathbf{P}_{i}$. For the second bad event to happen, letting $I^{\prime}\mathrel{\mathop{\ordinarycolon}}=\{i\in[L]\mathrel{\mathop{\ordinarycolon}}P^{*}_{i}\neq\emptyset\}$ (so $I^{\prime}\subseteq I$, $|I^{\prime}|\leq q$ and $|I|$ $\leq 2q$), there exists either (1) an $i\in I^{\prime}$ with $P_{i}^{*}=\{x^{*}\}$ such that $\bm{\rho}_{i}(x^{*})\neq\bm{\rho}_{i}(y)$ for some $y\in P_{i}$, or (2) an $i\in I\setminus I^{\prime}$ such that $\bm{\rho}_{i}(x)\neq\bm{\rho}_{i}(y)$ for some $x,y\in P_{i}$. The probability of the former, for each $i\in I^{\prime}$, is (given that we are in the yes case)
$1-\frac{|A_{i,1}|}{|j\in[n]\mathrel{\mathop{\ordinarycolon}}x_{j}^{*}=1|}\leq O\left(\frac{|P_{i}|\log n}{n}\right)$
if $\rho_{i}^{*}(x^{*})=1$ or
$1-\frac{|A_{i,0}|}{|j\in[n]\mathrel{\mathop{\ordinarycolon}}x_{j}^{*}=0|}\leq O\left(\frac{|P_{i}|\log n}{n}\right)$
if $\rho_{i}^{*}(x^{*})=0$. The probability of the latter, for some $i\in I\setminus I^{\prime}$, is at most
$1-\frac{|A_{i,0}|+|A_{i,1}|}{n}\leq O\left(\frac{|P_{i}|\log n}{n}\right).$
So conditioning on the first bad event not happening, the second bad event happens with probability at most
$\sum_{i\in I}O\left(\frac{|P_{i}|\log n}{n}\right)\leq\alpha,$
by using $\sum_{i\in I}|P_{i}|\leq 2q=2n/(C_{0}\log n)$ and setting $C_{0}$ to be sufficiently large. ∎
Finally we prove 27:
Proof of 27.
Fix a good $U=(Q,S,(P;R;\rho))$ in the support of $\mathcal{H}_{\text{yes}}$. Let $I\mathrel{\mathop{\ordinarycolon}}=\{i\in[L]\mathrel{\mathop{\ordinarycolon}}P_{i}\neq\emptyset\}$. For each $i\in I$, let $A_{i,1}$ and $A_{i,0}$ be defined as in the proof of 30. Given that $U$ is good, $|A_{i,0}|$ and $|A_{i,1}|$ satisfy the same inequalities established in the proof of 30.
Consider any fixed $T$ such that the probability of $(P;R;\rho)=\mathsf{outcome}(f_{T,{\bm{H}}},Q\cup S)$ is positive when ${\bm{H}}\sim\mathcal{E}_{\text{yes}}$. Then it suffices to show that
$\dfrac{\operatorname{{\bf Pr}}_{{\bm{H}}\sim\mathcal{E}_{\text{no}}}[\mathsf{outcome}(f_{T{\bm{H}}},Q\cup S)=(P;R;\rho)]}{\operatorname{{\bf Pr}}_{{\bm{H}}\sim\mathcal{E}_{\text{yes}}}[\mathsf{outcome}(f_{T,{\bm{H}}},Q\cup S)=(P;R;\rho)]}\geq\frac{1}{1+\alpha}.$
By definition of the distributions $\mathcal{E}_{\text{no}}$ and $\mathcal{E}_{\text{yes}}$, this ratio is a product over $i\in I$.
For each $i\in I$ with $\rho_{i}(x)=1$ for $x\in P_{i}$, the factor is
$\frac{\frac{2}{3}\cdot\frac{|A_{i,0}|}{n}+\frac{1}{3}}{\frac{2}{3}\cdot\frac{|A_{i,1}|}{n}}=\frac{(2/3)\cdot|A_{i,0}|+(n/3)}{(2/3)\cdot|A_{i,1}|}.$ (8)
Given that $|A_{i,1}|=(3n/4)\pm O(|P_{i}|\log n)$ and $|A_{i,0}|=(n/4)\pm O(|P_{i}|\log n)$ (and $|P_{i}|\log n\leq 2q\log n$ which can be set to be at most a sufficiently small constant multiple of $n$ by setting $C_{0}$ to be sufficiently small),
$\frac{(2/3)\cdot|A_{i,0}|+(n/3)}{(2/3)\cdot|A_{i,1}|}=\frac{(n/2)\pm O(|P_{i}|\log n)}{(n/2)\pm O(|P_{i}|\log n)}=1\pm O\left(\frac{|P_{i}|\log n}{n}\right).$ (9)
For each $i\in I$ with $\rho_{i}(x)=0$ for $x\in P_{i}$, the factor is
$\frac{\frac{2}{3}\cdot\frac{|A_{i,1}|}{n}}{\frac{2}{3}\cdot\frac{|A_{i,0}|}{n}+\frac{1}{3}}=\frac{(2/3)\cdot A_{i,1}}{(2/3)\cdot A_{i,0}+(n/3)}=1\pm O\left(\frac{|P_{i}|\log n}{n}\right).$ (10)
As a result, using $\sum_{i\in I}|P_{i}|\leq 2q=2n/(C_{0}\log n)$ and setting $C_{0}$ to be sufficiently large, we have
$\dfrac{\operatorname{{\bf Pr}}_{{\bm{H}}\sim\mathcal{E}_{\text{no}}}[\mathsf{outcome}(f_{T,{\bm{H}}},Q\cup S)=(P;R;\rho)]}{\operatorname{{\bf Pr}}_{{\bm{H}}\sim\mathcal{E}_{\text{yes}}}[\mathsf{outcome}(f_{T,{\bm{H}}},Q\cup S)=(P;R;\rho)]}$ $\geq\prod_{i\in I}\left(1-O\left(\dfrac{|P_{i}|\log n}{n}\right)\right)\geq\frac{1}{1+\alpha}.$
@@ 6 A two-sided adaptive lower bound
Observations similar to those below the statement of Theorem 14 show that Theorem 31 implies Theorem 3. As it will become clear in Section 6.1, all functions used in our lower bound proof are two-layer functions as well. It follows from 15 that it suffices to show that any randomized, adaptive, $\mathrm{MQ}$-only algorithm for testing two-layer functions with success probability $(99/100)\cdot 2/3$ must make $\tilde{\Omega}((\log n)^{2/3})$ queries. So in the rest of the section, we focus on $\mathrm{MQ}$-only algorithms.
@@ 6.1 Distributions $\mathcal{D}_{\mathrm{yes}}$ and $\mathcal{D}_{\mathrm{no}}$
    Sample ${\bm{H}}=(\bm{h}_{i,j}\mathrel{\mathop{\ordinarycolon}}i\in[L],j\in[M])$ from $\mathcal{E}_{\text{yes}}$, where each ${\bm{h}}_{i,j}\mathrel{\mathop{\ordinarycolon}}\{0,1\}^{n}\rightarrow\{0,1\}$ is (1) with probability $2/3$, a random dictatorship Boolean function, i.e., ${\bm{h}}_{i}(x)=x_{k}$ with $k$ sampled uniformly at random from $[n]$; and (2) with probability $1/3$, the all-0 function.
@@
Note that the number of points on the $3n/4$ layer is ${n\choose 3n/4}\geq|f^{-1}(1)|/100$. Thus, the expected size of $\bm{X}$ is at least $|f^{-1}(1)|/10000$, and $|\bm{X}|=\Omega(|f^{-1}(1)|)$ with probability at least 0.0001. ∎